# Toward Defining an Efficient and Expandable File Format for AI-Generated Contents

### v1

@@ I Introduction
Notably, most commonly used image file formats like PNG [7] and JPEG-XL [8, 9] compress an image into a file by performing the transform and entropy coding to reduce the spatial redundancy. Despite that, image compression in the pixel-wise space merely achieves the limited compression ratio of approximately 1/2 to 1/12 for high-fidelity image compression, which implies that it still requires hundreds of kilobytes or even several megabytes to store a high-resolution AI-generated image. To overcome the above challenge, we aim to break through the limitation of existing image file formats for AIGC images, achieving ultra-low bitrate compression while maintaining the high fidelity of AIGC images.
In this work, we define a new efficient and expandable AI-Generated Image Format, named AIGIF, which is capable of achieving high-fidelity AIGC image compression with an ultra-low bitrate. Notably, different from natural images, AIGC images are generated with generative models (e.g., Stable Diffusion v1.5) by user-defined data configuration like text prompts and size. Since the large computational cost, a particular platform (e.g., CPU or GPU) is required to support the generation process. Meanwhile, the creation process and generation syntax of AIGC images are usually accessible for users, which raises an intuitive question for us: “whether we can compress the AIGC images by compressing generation syntax only instead of images themselves like existing image file formats”.
@@ II-A Preliminary
AIGC has paved the path to general artificial intelligence(AGI) by empowering AI models with the powerful creation, understanding, and interactive capabilities. As the representative task of AIGC, text-to-image generation has progressed significantly with the advancement of diffusion models [12, 13, 14] and multi-modality methods like CLIP [15]. In this work, we adopt the popular Stable Diffusion [1] to produce the AIGC images. Given the fine-grained description prompts, users can generate amounts of high-quality images they want, which can assist the artistic creation or entertainment, posing a significant challenge to existing image file formats in terms of storage and transmission.
@@ II-B4 Overall Pipeline
Additionally, to handle the loss of generation syntax elements in special cases, our AIGIF supports direct compression and transmission of original AIGC images (shown as dashed lines in Fig. 1). This feature ensures high-fidelity image preservation and transmission even when the generation syntax cannot be fully utilized. By employing a well-defined composable bitstream structure and offering flexible options for image compression and recreation, AIGIF provides a solution for efficiently reproducing high-quality AI-generated images.
@@ III-B Results
TABLE II: File size comparison of different image formats.
@@
Complexity. As shown in Table III, the time complexity of our AIGIF comprises two main components: the image generation process and the entropy coding of generation information. Notably, AIGIF involves a user-defined generative process, making the total time complexity highly dependent on the image generation time. Although the standard SD1.5 model typically takes around 50 seconds to generate a 1024$\times$1024 image, considerable research has focused on accelerating the diffusion process. Techniques such as distillation [19, 20, 21], consistency models [22, 23], and faster samplers [24, 25] have significantly reduced the required diffusion steps. These acceleration techniques are orthogonal to our approach. Here we provide the time complexity of two models, SD1.5 and Hyper-SD, where Hyper-SD is a SOTA diffusion model acceleration technique [21]. As demonstrated in Table III, the 1-step Hyper-SD method only requires about one second for image generation, significantly reducing decoding complexity.
@@ IV Limitations
The potential limitation of our AIGIF is the computational cost on the decoder side, which arises from generation process and is the same as the challenge faced by existing diffusion model-based image compression methods [27, 28, 29, 30, 31]. We believe it can be effectively eliminated with the rapid development of efficient image generation methods [32, 33, 34, 35, 36, 37].

### v3

# Towards Defining an Efficient and Expandable File Format for AI-Generated Contents

@@ I Introduction
Notably, most commonly used image file formats like PNG [7] and JPEG-XL [8, 9] compress an image into a file by performing the transform and entropy coding to reduce the spatial redundancy. Despite that, image compression in the pixel-wise space merely achieves the limited compression ratio of approximately 1/2 to 1/12 for high-fidelity image compression [10, 11], which implies that it still requires hundreds of kilobytes or even several megabytes to store a high-resolution AI-generated image. To overcome the above challenge, we aim to break through the limitation of existing image file formats for AIGC images, achieving ultra-low bitrate compression while maintaining the high fidelity of AIGC images.
In this work, we define a new efficient and expandable AI-Generated Image Format, named AIGIF, which is capable of achieving high-fidelity AIGC image compression with an ultra-low bitrate. Notably, unlike natural images, AIGC images are generated with generative models (e.g., Stable Diffusion v1.5) by user-defined data configuration like text prompts and size. Since the large computational cost, a particular platform (e.g., CPU or GPU) is required to support the generation process. Meanwhile, the creation process and generation syntax of AIGC images are usually accessible for users, which raises an intuitive question for us: “whether we can compress AIGC images by compressing generation syntax only instead of images themselves like existing image file formats”.
@@ II-A Preliminary
AIGC has paved the path to general artificial intelligence(AGI) by empowering AI models with the powerful creation, understanding, and interactive capabilities. As the representative task of AIGC, text-to-image generation has progressed significantly with the advancement of diffusion models [14, 15, 16, 17] and multi-modality methods like CLIP [18]. In this work, we adopt the popular Stable Diffusion [1] to produce the AIGC images. Given the fine-grained description prompts, users can generate amounts of high-quality images they want, which can assist the artistic creation or entertainment, posing a significant challenge to existing image file formats in terms of storage and transmission.
@@ II-B4 Overall Pipeline
Additionally, to handle the loss of generation syntax elements in special cases, our AIGIF supports direct compression and transmission of original AIGC images (shown as dashed lines in Fig. 1) [7, 8, 19, 20, 21]. This feature ensures high-fidelity image preservation and transmission even when the generation syntax cannot be fully utilized. By employing a well-defined composable bitstream structure and offering flexible options for image compression and recreation, AIGIF provides a solution for efficiently reproducing high-quality AI-generated images.
@@ III-B Results
TABLE II: Comparisons of different image file formats on AIGC images.
@@
Complexity. We present decoding time comparisons in Table II and a detailed complexity analysis in Table III. As shown in Table III, the time complexity of our AIGIF comprises two main components: the image generation process and the entropy coding of generation information. Notably, AIGIF involves a user-defined generative process, making the total time complexity highly dependent on the image generation time. Although the standard SD1.5 model typically takes around 50 seconds to generate a 1024$\times$1024 image, considerable research has focused on accelerating the diffusion process. Techniques such as distillation [25, 26, 27], consistency models [28, 29], and faster samplers [30, 31] have significantly reduced the required diffusion steps. These acceleration techniques are orthogonal to our approach. Here we provide the time complexity of two models, SD1.5 and Hyper-SD, where Hyper-SD is a SOTA diffusion model acceleration technique [27]. As demonstrated in Table III, the 1-step Hyper-SD method only requires about one second for image generation, significantly reducing decoding complexity.
@@ IV Limitations
The potential limitation of our AIGIF is the computational cost on the decoder side, which arises from generation process and is the same as the challenge faced by existing diffusion model-based image compression methods [33, 34, 35, 36, 37, 38]. We believe it can be effectively eliminated with the rapid development of efficient image generation methods [39, 40, 41, 42, 43].